# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

### v1

@@ 4 Robust Bellman Operator
The operator $\mathbf{T}_{g}$ transforms a candidate value function $V$ by incorporating the worst-case transition effect. A key challenge in solving the robust Bellman equation is that $\mathbf{T}_{g}$ does not satisfy contraction under standard norms, preventing the use of conventional fixed-point iteration. To cope with this problem, we establish that $\mathbf{T}_{g}$ is a contraction under the span semi-norm. This allows us to develop provably efficient stochastic approximation algorithms. Throughout this paper, we make the following standard assumption regarding the structure of the induced Markov chain.
@@
Assumption 1 is used widely in all robust average reinforcement learning literatures (Wang et al., 2023a, b, 2024; Sun et al., 2024). This assumption ensures that, under any transition model within the uncertainty set, the policy $\pi$ induces a single recurrent communicating class. A well-known result in average-reward MDPs states that under Assumption 1, the average reward is independent of the starting state, i.e., for any $\mathsf{P}\in\mathcal{P}$ and all $s,s^{\prime}\in\mathcal{S}$,
$g^{\pi}_{\mathsf{P}}(s)=g^{\pi}_{\mathsf{P}}(s^{\prime}).$ (14)
Thus, we can drop the dependence on the initial state and simply write $g^{\pi}_{\mathsf{P}}$ as the robust average reward.
Under Assumption 1, we are able to establish the semi-norm contraction property. Before proceeding, we first establish the semi-norm property of non-robust average reward bellman operator for a policy $\pi$ under transition $\mathsf{P}$ defined as follows.
@@
Let $\mathcal{S}$ be a finite state space, and let $\pi$ be a stationary policy. If the Markov chain induced by $\pi$ under the transition $\mathsf{P}$ is irreducible and aperiodic, there exists a constant $\beta\in(0,1)$ such that for all $V_{1},V_{2}\in\mathbb{R}^{S}$ and $g\in\mathbb{R}$
$\|\mathbf{T}_{g}^{\mathsf{P}}(V_{1})-\mathbf{T}_{g}^{\mathsf{P}}(V_{2})\|_{\mathrm{sp}}\leq\beta\|V_{1}-V_{2}\|_{\mathrm{sp}},$ (16)
where
$\|v\|_{\mathrm{sp}}\coloneqq\max_{s}\,v(s)-\min_{s}\,v(s).$
The proof of Lemma 1 is in Appendix A.1, where the properties of irreduible and aperiodic finite state Markov chain is utilized. Thus we show the (non-robust) average reward bellman operator $\mathbf{T}_{g}^{\mathsf{P}}$ is a strict contraction under the span semi-norm. Based on the above results, we now formally establish the contraction property of the robust average reward bellman operator by leveraging Lemma 1 and the compactness of the uncertainty sets.
@@
Under Assumption 1, and if $\mathcal{P}$ is compact, the robust bellman operator $\mathbf{T}_{g}$ is a contractive mapping with respect to the span semi-norm for any $g$. Specifically, there exist $\gamma\in(0,1)$ such that
$\|\mathbf{T}_{g}(V_{1})-\mathbf{T}_{g}(V_{2})\|_{\mathrm{sp}}\leq\gamma\|V_{1}-V_{2}\|_{\mathrm{sp}},\quad\forall V_{1},V_{2}\in\mathbb{R}^{S},g\in\mathbb{R}$ (17)
where
$\|v\|_{\mathrm{sp}}\coloneqq\max_{s}\,v(s)-\min_{s}\,v(s).$
@@
The proof of Theorem 2 is in Appendix A.2. Since all the uncertainty sets listed in Section 3.2 are closed and bounded in a real vector space, these uncertainty sets are all compact and satisfy the comtraction property in Theorem 2.
5 Convergence of Span Contraction with Bias
In the previous section, we established that the robust Bellman operator is a contraction under the span semi-norm, ensuring that policy evaluation can be analyzed within a well-posed stochastic approximation framework. However, conventional stochastic approximation methods typically assume unbiased noise, where variance diminishes over time without introducing systematic drift. In contrast, the noise in robust policy evaluation under TV and Wasserstein distance uncertainty sets exhibits a small but persistent bias, arising from the estimators of the support functions $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ (discussed in Section 6). This bias, if not properly addressed, can lead to uncontrolled error accumulation, affecting the reliability of policy evaluation. To address this challenge, this section introduces a novel analysis of biased stochastic approximation, leveraging properties of dual norms to ensure that the bias remains controlled and does not significantly impact the convergence rate. Our results extend prior work on unbiased settings and provide the first explicit finite-time guarantees, which is further used to establish the sample complexity of policy evaluation in robust average reward RL. Specifically, we analyze the iteration complexity for solving the fixed equivalent class equation $H(x^{*})-x^{*}\in\overline{E}$ where $\overline{E}\coloneqq\{c\mathbf{e}:c\in\mathbb{R}\}$ with $\mathbf{e}$ being the all-ones vector. The stochastic approximation iteration being used is as follows:
with $\eta_{t}>0$ being the step-size sequance and with the following assumptions on the operator $H$ and noise $\omega^{t}$:
  • $H$ is a contractive mapping with respect to the span semi-norm, there exist $\gamma\in(0,1)$ such that
    $\|H(x)-H(y)\|_{\mathrm{sp}}\leq\gamma\,\|x-y\|_{\mathrm{sp}},\quad\forall x,y$ (19)
  • the noise terms $\omega^{t}$ are i.i.d. and have bounded bias and variance
    $\mathbb{E}[\,\|w^{t}\|_{\mathrm{sp}}^{2}|\mathcal{F}^{t}]\leq A+B\,\|x^{t}-x^{*}\|_{\mathrm{sp}}^{2}\quad\text{and}\quad\bigl{\|}\mathbb{E}[\,w^{t}|\mathcal{F}^{t}]\bigr{\|}_{\mathrm{sp}}\leq\varepsilon_{\mathrm{bias}}$ (20)
Theorem 3.
$\mathbb{E}\Bigl{[}\|x^{T}-x^{*}\|^{2}_{\mathrm{sp}}\Bigr{]}\leq\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}\right)\|x^{0}-x^{*}\|^{2}_{\mathrm{sp}}+\operatorname*{\mathcal{O}}\left(\frac{A}{(1-\gamma)^{2}T}\right)+\operatorname*{\mathcal{O}}\left(\frac{x_{\mathrm{sp}}\varepsilon_{\text{bias}}\log T}{1-\gamma}\right)$ (21)
where $x_{\mathrm{sp}}\coloneqq\sup_{x}\|x\|_{\mathrm{sp}}$ is the upper bound of the span for all $x^{t}$.
Theorem 3 adapts the analysis of (Zhang et al., 2021) and extends it to a biased i.i.d. noise setting. To manage the bias terms, we leverage properties of dual norms (see (59)-(66) in Appendix B.3) to bound the inner product between the error term and the gradient, ensuring that the bias influence remains logarithmic in $T$ rather than growing unbounded, while also carefully structuring the stepsize decay to mitigate long-term accumulation. This results in an extra $\varepsilon_{\mathrm{bias}}$ term with logarithmic dependence of the total iteration $T$. The detailed proof of Theorem 3 along with the exact constant terms is in Appendix B.
6 Queries from Uncertainty Set
@@
6.1 Linear Contamination Uncertainty Set
Recall that the $\delta$-contamination uncertainty set is $\mathcal{P}^{a}_{s}=\{(1-\delta)\mathsf{P}^{a}_{s}+\delta q:q\in\Delta(\mathcal{S})\},$ where $10<\delta<$ 1 is the radius. Since the support function can be computed by (9) and the expression is linear in the nominal transition kernel $\mathsf{P}^{a}_{s}$. A direct approach is to use the transition to the subsequent state to construct our estimator:
@@ 6.1 Linear Contamination Uncertainty Set
where $s^{\prime}$ is a subsequent state sample after $(s,a)$. Hence, the sample complexity of (22) is just one. A well know result from (Wang et al., 2023b) is that $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained by (22) is unbiased and has bounded variance as follows:
@@
6.2 Non-Linear Uncertainty Sets
Non-linear uncertainty sets such as TV distance uncertainty set and Wasserstein distance uncertainty sets have a non-linear relationship between the nonminal distribution $\mathsf{P}^{a}_{s}$ and the support function ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$. Previous works such as (Blanchet and Glynn, 2015; Blanchet et al., 2019; Wang et al., 2023b) have proposed a multi-level Monte-Carlo (MLMC) method for obtaining an unbiased estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ with bounded variance. However, their approach all require drawing $2^{N+1}$ samples where $N$ is sampled from a geometric distribution $\mathrm{Geom}(\Psi)$ with parameter $\Psi\in(0,0.5)$. This operation would need infinite samples in expectation for obtaining each single estimator:
$\mathbb{E}[2^{N+1}]=\sum^{\infty}_{N=0}2^{N+1}\Psi(1-\Psi)^{N}=\sum^{\infty}_{N=0}2\Psi(2-2\Psi)^{N}\rightarrow\infty$ (24)
To handle the above problem, we aim to provide an estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ with finite sample complexity and small enough bias. We construct a level-MLMC estimator under geometric sampling with parameter $\Psi=0.5$ as shown in Algorithm 1.
Algorithm 1 Truncated MLMC Estimator for TV and Wasserstein Unceretainty Sets
Input: $s\in\mathcal{S}$, $a\in\mathcal{A}$, Truncation level $N_{\max}$, Value function $V$
@@ 6.2 Non-Linear Uncertainty Sets
3:Collect $2^{N^{\prime}+1}$ i.i.d. samples of $\{s^{\prime}_{i}\}^{2^{N^{\prime}+1}}_{i=1}$ with $s^{\prime}_{i}\sim\mathsf{P}^{a}_{s}$ for each $i$
@@
8:if TV distance uncertainty set then Obtain $\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}}(V)$ from (10)
9:else if Wasserstein distance uncertainty set then Obtain $\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}}(V)$ from (12)
@@
6.2.1 Sample Complexity for Querying Non-Linear Uncertainty Sets
@@
The proof of Theorem 4 is in Appendix C.1, which demonstrates that setting the geometric sampling parameter to $\Psi=0.5$ ensures that the expected number of samples follows a linear growth pattern rather than an exponential one. This choice precisely cancels out the effect of the exponential sampling inherent in the truncated MLMC estimator, preventing infinite expected sample complexity. This result shows that the expected number of queries grows only linearly with $N_{\max}$, ensuring that the sampling cost remains manageable even for large truncation levels. The key factor enabling this behavior is setting the geometric distribution parameter to $0.5,$ which balances the probability mass across different truncation levels, preventing an exponential increase in sample complexity.
6.2.2 Exponential Bias Decay
@@
and under Wasserstein uncertainty set, we have:
@@
The proof of Theorem 5 is in Appendix C.2. One important lemma used in the proof is the following Lemma 2, where we show the Lipschitz property for both TV and Wasserstein distance uncertainty sets.
@@
$|\sigma_{\mathcal{P}_{TV}}(V)-\sigma_{\mathcal{Q}_{TV}}(V)|\leq(1+\frac{1}{\delta})\|V\|_{\mathrm{sp}}\|p-q\|_{1}$ (28)
$|\sigma_{\mathcal{P}_{W}}(V)-\sigma_{\mathcal{Q}_{W}}(V)|\leq\|V\|_{\mathrm{sp}}\|p-q\|_{1}$ (29)
We refer the proof of Theorem 5 to Appendix C.3.
6.2.3 Linear Variance
@@
Theorem 6 establishes that the variance of the truncated MLMC estimator grows linearly with $N_{\max}$, ensuring that the estimator remains stable even as the truncation level increases. The proof of Theorem 6 is in Appendix C.4, which follows from bounding the second moment of the estimator by analyzing the variance decomposition across different MLMC levels. Specifically, by expressing the estimator in terms of successive refinements of the transition model, we show that the variance accumulates additively across levels due to the binomial concentration property.
7 Propoosed Algorithm and Final Results
We present the formal algorithm for robust policy evaluation and robust average reward for a given policy $\pi$ in Algorithm 2. Algorithm 2 presents a robust temporal difference (TD) learning method for policy evaluation in robust average-reward MDPs. This algorithm builds upon the truncated MLMC estimator (Algorithm 1) and the biased stochastic approximation framework in Section 5, ensuring both efficient sample complexity and finite-time convergence guarantees.

### v2

@@ 2.2 Robust Bellman Operator
The operator $\mathbf{T}_{g}$ transforms a value function $V$ by incorporating the worst-case transition effect. A key challenge in solving the robust Bellman equation is that $\mathbf{T}_{g}$ does not satisfy contraction under standard norms, preventing the use of conventional fixed-point iteration. To cope with this problem, we establish that $\mathbf{T}_{g}$ is a contraction under some constructed semi-norm. This allows us to develop provably efficient stochastic approximation algorithms. Throughout this paper, we make the following standard assumption regarding the structure of the induced Markov chain.
@@
Assumption 1 is used widely in the current robust average reinforcement learning literature Wang et al. (2023a, b, 2024); Sun et al. (2024). This assumption ensures that, under any transition model within the uncertainty set, the policy $\pi$ induces a single recurrent communicating class. A well-known result in average-reward MDPs states that under Assumption 1, the average reward is independent of the starting state, i.e., for any $\mathsf{P}\in\mathcal{P}$ and all $s,s^{\prime}\in\mathcal{S}$, we have $g^{\pi}_{\mathsf{P}}(s)=g^{\pi}_{\mathsf{P}}(s^{\prime})$. Thus, we can drop the dependence on the initial state and simply write $g^{\pi}_{\mathsf{P}}$ as the robust average reward.
Remark 1.
We note that when the radius of uncertainty sets are small enough, Assumption 1 could be relaxed into assuming only the nominal transition being irreducible and aperiodic. We provide the detailed discussions in Appendix B.
3 Semi-Norm Contraction of Robust Bellman Operators
Under Assumption 1, we are able to establish the semi-norm contraction property. For motivation, we first establish the semi-norm contraction property of the non-robust average reward Bellman operator for a policy $\pi$ under transition $\mathsf{P}$ defined as follows:
@@ 3 Semi-Norm Contraction of Robust Bellman Operators
Let $\mathcal{S}$ be a finite state space, and let $\pi$ be a stationary policy. If the Markov chain induced by $\pi$ under the transition $\mathsf{P}$ is irreducible and aperiodic, there exists a semi-norm $\|\cdot\|_{\mathsf{P}}$ with kernel $\{c\mathbf{e}:c\in\mathbb{R}\}$ and a constant $\beta\in(0,1)$ such that for all $V_{1},V_{2}\in\mathbb{R}^{S}$ and any $g\in\mathbb{R}$,
$\bigl{\|}\mathbf{T}_{g}^{\mathsf{P}}(V_{1})-\mathbf{T}_{g}^{\mathsf{P}}(V_{2})\bigr{\|}_{\mathsf{P}}\leq\beta\|V_{1}-V_{2}\|_{\mathsf{P}}.$ (12)
Proof Sketch
Under ergodicity, the one‐step transition matrix (denoted as $\mathsf{P}^{\pi}$) has a unique stationary distribution $d^{\pi}$, define the stationary projector $E=\mathbf{e}^{\top}d^{\pi}$, then the fluctuation matrix (defined as $Q^{\pi}=\mathsf{P}^{\pi}-E$) has all eigenvalues strictly inside the unit circle. Standard finite‐dimensional theory (via the discrete Lyapunov equation Horn and Johnson (2012)) would produce a norm $\|\cdot\|_{Q}$ on $\mathbb{R}^{S}$ such that there is a constant $\alpha\in(0,1)$ such that for any $x\in\mathbb{R}^{S}$, $\|Q^{\pi}x\|_{Q}\leq\alpha\|x\|_{Q}$. We then build the semi-norm as follows:
$\|x\|_{\mathsf{P}}=\|Q^{\pi}x\|_{Q}+\epsilon\inf_{c\in\mathbb{R}}\|x-c\mathbf{e}\|_{Q},\quad 0<\epsilon<1-\alpha,$ (13)
so that its kernel is exactly the constant vectors (the second term vanishes only on shifts of $\mathbf{e}$) and the first term enforces a one‐step shrinkage by $\beta=\alpha+\epsilon<1$. A short calculation then shows $\|\mathsf{P}^{\pi}x\|_{\mathsf{P}}\leq\beta\|x\|_{\mathsf{P}}$, yielding the desired contraction, which leads to the overall result.
The concrete proof of Lemma 1 including the detailed construction of the semi-norm $\|\cdot\|_{\mathsf{P}}$ is in Appendix C.1, where the properties of irreducible and aperiodic finite state Markov chain are utilized. Thus, we show the (non-robust) average reward Bellman operator $\mathbf{T}_{g}^{\mathsf{P}}$ is a strict contraction under $\|\cdot\|_{\mathsf{P}}$. Based on the above motivations, we now formally establish the contraction property of the robust average reward Bellman operator by leveraging Lemma 1 and the compactness of the uncertainty sets.
@@
Under Assumption 1, and if $\mathcal{P}$ is compact, there exists a semi-norm $\|\cdot\|_{\mathcal{P}}$ with kernel $\{c\mathbf{e}:c\in\mathbb{R}\}$ such that the robust Bellman operator $\mathbf{T}_{g}$ is a contraction. Specifically, there exist $\gamma\in(0,1)$ such that
$\|\mathbf{T}_{g}(V_{1})-\mathbf{T}_{g}(V_{2})\|_{\mathcal{P}}\;\leq\;\gamma\,\|V_{1}-V_{2}\|_{\mathcal{P}},\;\forall\;V_{1},V_{2}\in\mathbb{R}^{S},\,g\in\mathbb{R}.$ (14)
Proof Sketch
For any $\mathsf{P}\in\mathcal{P}$, the one‐step transition matrix $\mathsf{P}^{\pi}$ has a unique stationary projector $E_{\mathsf{P}}$ due to ergodicity. Since $\mathcal{P}$ is compact, the family of fluctuation matrices $\{Q_{\mathsf{P}}^{\pi}=\mathsf{P}^{\pi}-E_{\mathsf{P}}:\mathsf{P}\in\mathcal{P}\}$ has joint spectral radius strictly less than $1$. By Lemma 10in Berger and Wang (1992), one is able to construct an “extremal norm” (denoted as $\|\cdot\|_{\rm ext}$) under which every $Q_{\mathsf{P}}^{\pi}$ contracts by a uniform factor $\alpha\in(0,1)$. Mimicking the non-robust case in Lemma 1, we similarly define
$\|x\|_{\mathcal{P}}=\sup_{\mathsf{P}\in\mathcal{P}}\|Q_{\mathsf{P}}^{\pi}x\|_{\rm ext}+\epsilon\inf_{c\in\mathbb{R}}\|x-c\mathbf{e}\|_{\rm ext},\quad 0<\epsilon<1-\alpha.$ (15)
The supremum term zeros out if $x\in\{c\mathbf{e}:c\in\mathbb{R}\}$ , and it inherits the uniform one‐step shrinkage by $\alpha$. Adding the small quotient term fixes the kernel without spoiling $\gamma=\alpha+\epsilon<1$, so one shows at once
$\|\mathbf{T}^{\mathsf{P}}_{g}(V_{1})-\mathbf{T}^{\mathsf{P}}_{g}(V_{2})\|_{\mathcal{P}}\leq\gamma\|V_{1}-V_{2}\|_{\mathcal{P}}\quad\text{for all}\;\mathsf{P}\in\mathcal{P}$ (16)
The above leads to the desired results.
@@
The concrete proof of Theorem 2 along with the detailed construction of the semi-norm $\|\cdot\|_{\mathcal{P}}$ are in Appendix C.2. Since all the uncertainty sets listed in Section 2.1 are closed and bounded in a real vector space, these uncertainty sets are all compact and satisfy the contraction property in Theorem 2.
4 Efficient Estimators for Uncertainty Sets
To utilize the contraction property in Section 3 to obtain convergence rate results, our idea is perform the following iterative stochastic approximation:
$V_{t+1}(s)\leftarrow V_{t}(s)+\eta_{t}\left(\hat{\mathbf{T}}_{g}(V_{t})(s)-V_{t}(s)\right),\quad\forall s\in\mathcal{S}$ (17)
where the learning rate $\eta_{t}$ would be specified in Section 5. The detailed analysis and complexities of the general stochastic approximation in the form of (17) is provided in Appendix D. Theorem 8 implies that if $\hat{\mathbf{T}}_{g}(V)$, being an estimator of $\mathbf{T}_{g}(V)$, could be constructed with bounded variance and small bias, $V_{t}$ converges to a solution of the Bellman equation in (4). However, the challenge of constructing our desired $\hat{\mathbf{T}}_{g}(V)$ lies in the construction of the support function estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$.
@@ 4 Efficient Estimators for Uncertainty Sets
Linear Contamination Uncertainty Set
Recall that the $\delta$-contamination uncertainty set is $\mathcal{P}^{a}_{s}=\{(1-\delta)\tilde{\mathsf{P}}^{a}_{s}+\delta q:q\in\Delta(\mathcal{S})\},$ where $10<\delta<$ 1 is the radius. Since the support function can be computed by (6) and the expression is linear in the nominal transition kernel $\tilde{\mathsf{P}}^{a}_{s}$. A direct approach is to use the transition to the subsequent state to construct our estimator:
@@
where $s^{\prime}$ is a subsequent state sample after $(s,a)$. Hence, the sample complexity of (18) is just one. Lemma 12 from Wang et al. (2023b) states that $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ obtained by (18) is unbiased and has bounded variance as follows:
@@
Nonlinear Contamination Sets
Regarding TV and Wasserstein distance uncertainty sets, they have a nonlinear relationship between the nominal distribution $\tilde{\mathsf{P}}^{a}_{s}$ and the support function ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$. Previous works such as (Blanchet and Glynn, 2015; Blanchet et al., 2019; Wang et al., 2023b) have proposed a multi-level Monte-Carlo (MLMC) method for obtaining an unbiased estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$ with bounded variance. However, their approaches require drawing $2^{N+1}$ samples where $N$ is sampled from a geometric distribution $\mathrm{Geom}(\Psi)$ with parameter $\Psi\in(0,0.5)$. This operation would need infinite samples in expectation for obtaining each single estimator as $\mathbb{E}[2^{N+1}]=\sum^{\infty}_{N=0}2^{N+1}\Psi(1-\Psi)^{N}=\sum^{\infty}_{N=0}2\Psi(2-2\Psi)^{N}\rightarrow\infty$. To handle the above problem, we aim to provide an estimator $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ with finite sample complexity and small enough bias. We construct a truncated-MLMC estimator under geometric sampling with parameter $\Psi=0.5$ as shown in Algorithm 1.
Algorithm 1 Truncated MLMC Estimator for TV and Wasserstein Uncertainty Sets
Input: $s\in\mathcal{S}$, $a\in\mathcal{A}$, Max level $N_{\max}$, Value function $V$
@@
3:Collect $2^{N^{\prime}+1}$ i.i.d. samples of $\{s^{\prime}_{i}\}^{2^{N^{\prime}+1}}_{i=1}$ with $s^{\prime}_{i}\sim\tilde{\mathsf{P}}^{a}_{s}$ for each $i$
@@
8:if TV then Obtain $\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}}(V)$ from (7)
9:else if Wasserstein then Obtain $\sigma_{\hat{\mathsf{P}}^{a,1}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,E}_{s,N^{\prime}+1}}(V),\sigma_{\hat{\mathsf{P}}^{a,O}_{s,N^{\prime}+1}}(V)$ from (9)
@@
The proof of Theorem 3 is in Appendix E.1, which demonstrates that setting the geometric sampling parameter to $\Psi=0.5$ ensures that the expected number of samples follows a linear growth pattern rather than an exponential one. This choice precisely cancels out the effect of the exponential sampling inherent in the truncated MLMC estimator, preventing infinite expected sample complexity. This result shows that the expected number of queries grows only linearly with $N_{\max}$, ensuring that the sampling cost remains manageable even for large truncation levels. The key factor enabling this behavior is setting the geometric distribution parameter to $0.5,$ which balances the probability mass across different truncation levels, preventing an exponential increase in sample complexity.
@@
where $\delta$ denotes the radius of TV distance. Under Wasserstein uncertainty set, we have:
@@
The proof of Theorem 4 is in Appendix E.2. One important lemma used in the proof is the following Lemma 2, where we show the Lipschitz property for both TV and Wasserstein distance uncertainty sets.
@@
$|\sigma_{\mathcal{P}_{TV}}(V)-\sigma_{\mathcal{Q}_{TV}}(V)|\leq\|V\|_{\mathrm{sp}}\|p-q\|_{1}\;\text{and}\;|\sigma_{\mathcal{P}_{W}}(V)-\sigma_{\mathcal{Q}_{W}}(V)|\leq\|V\|_{\mathrm{sp}}\|p-q\|_{1}$ (23)
We refer the proof of Theorem 4 to Appendix E.3.
@@
Theorem 5 establishes that the variance of the truncated MLMC estimator grows linearly with $N_{\max}$, ensuring that the estimator remains stable even as the truncation level increases. The proof of Theorem 5 is in Appendix E.4, which follows from bounding the second moment of the estimator by analyzing the variance decomposition across different MLMC levels. Specifically, by expressing the estimator in terms of successive refinements of the transition model, we show that the variance accumulates additively across levels due to the binomial concentration property.
5 Robust Average Reward TD Learning
Equipped with the methods of constructing $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$, we now present the formal algorithm for robust policy evaluation and robust average reward for a given policy $\pi$ in Algorithm 2. Algorithm 2 presents a robust temporal difference (TD) learning method for policy evaluation in robust average-reward MDPs. This algorithm builds upon the truncated MLMC estimator (Algorithm 1) and the biased stochastic approximation framework in Section D, ensuring both efficient sample complexity and finite-time convergence guarantees.